OpenStudy (anonymous):

y=-3x^2-2x+1 Graph parabola,give its vertex,x and y intercept. Can someone help me solve this

OpenStudy (anonymous):

x interecept, set y=0, solve for x. y -intercept , set x =0 and solve for y.

OpenStudy (campbell_st):

well the vertex can be found using x = -b/2a which gives the line of symmetry for the parabola... and the vertex lies on the line of symmetry so in your question b = -2 and a = -3 so the line of symmetry is x = - 2/6 or x = -1/3 substitute this x value for find the corresponding y value.. when you'll have the vertex.. the y intercept is the easiest... set x = 0 and solve for y... the x- intercepts set y = 0 and solve for x and the quadratic will factor... or you can use the general quadratic formula to find the x intercepts. hope it helps..